### v2

@@ Abstract
Data-driven modeling of constrained multibody systems faces two persistent challenges: high computational cost and limited long-term prediction accuracy. To address these issues, we introduce the Flow-Matching Neural Ordinary Differential Equation (FNODE), a framework that learns acceleration vector fields directly from trajectory data. By reformulating the training objective to supervise accelerations rather than integrated states, FNODE eliminates the need for backpropagation through an ODE solver, which represents a bottleneck in traditional Neural ODEs. Acceleration targets are computed efficiently using numerical differentiation techniques, including a hybrid Fast Fourier Transform (FFT) and Finite Difference (FD) scheme. We evaluate FNODE on a diverse set of benchmarks, including the single and triple mass-spring-damper systems, double pendulum, slider-crank, and cart-pole. Across all cases, FNODE consistently outperforms existing approaches such as Multi-Body Dynamic Neural ODE (MBD-NODE), Long Short-Term Memory (LSTM) networks, and Fully Connected Neural Networks (FCNN), demonstrating good accuracy, generalization, and computational efficiency.
Keywords:
Multibody dynamics Neural ODE Constrained dynamics Scientific machine learning
MSC:
34C60 37M05
@@ 1 Introduction
Multibody dynamics simulation is relevant in engineering design and scientific computation, with applications in numerous fields, e.g., robotics featherstone2008rigid , vehicle and train dynamics shabana2020dynamics , biomechanics delp2007opensim , and molecular dynamics schlick2010molecular , to name a few. Traditionally, modeling multibody systems has relied on first-principles approaches such as Newton-Euler wittenburg2008dynamics , Lagrangian goldstein2002safko , or Hamiltonian mechanics marsden1999introduction . While these classical methods provide accurate and interpretable models, they are challenged in scenarios where the underlying dynamics are partially unknown, highly complex, or subject to unmodeled effects such as friction, contact forces, or material nonlinearities glocker2013set .
@@
Machine learning has opened the door to data-driven simulation of dynamical systems, offering the potential to learn complex behaviors directly from observational data raissi2019physics ; karniadakis2021physics . Early approaches in this domain employed classical machine learning architectures such as Long Short-Term Memory (LSTM) networks vlachas2018data ; hochreiter1997long and Fully Connected Neural Networks (FCNNs) sanchez2020learning to learn discrete-time mappings from current states to future states. While these methods have shown promise in short-term prediction tasks, they show limited out-of-distribution generalization. Specifically, these black-box predictors learn fixed-timestep transitions $f:s_{t}\mapsto s_{t+\Delta t}$, which not only limits their temporal resolution but also causes error accumulation that grows exponentially with the prediction horizon ren2024learning . Furthermore, such discrete-time approaches often fail to capture the continuous nature of physical dynamics or enforce fundamental conservation laws of conservative systemsjin2020sympnets .
Neural Ordinary Differential Equations (NODEs) chen2018neural marked a major step in continuous-time learning by representing system dynamics as neural vector fields integrated with standard ODE solvers. This formulation handles irregular time sampling and directly connects learning with established numerical integration theory kidger2020neural . This approach has been successfully applied to various physical systems, with extensions including Hamiltonian Neural Networks greydanus2019hamiltonian , Lagrangian Neural Networks cranmer2020discovering , and constrained formulations for systems with known invariantsyang2020learning . Despite their strengths, NODEs suffer from a bottleneck: the adjoint sensitivity method for backpropagation requires solving an additional backward ODE, which increases computational cost and memory use, and limits scalability for stiff or high-dimensional systems gholaminejad2019anode ; wang2024mbdnode .
@@
Recent advances in flow-matching algorithms liu2023flow ; lipman2023flow show that vector fields can be learned by directly supervising derivatives at sampled points, avoiding backward ODE solves tong2024improving . Applying this idea to dynamical systems, we train neural networks to approximate accelerations directly, eliminating costly integration in the learning loop li2020fourier . We propose Flow-Matching Neural Ordinary Differential Equations (FNODE), a framework for efficient data-driven modeling of multibody systems. Instead of predicting future states via time integration, FNODE trains a neural network to learn directly the acceleration vector field – the natural quantity in mechanical systems governed by force laws arnol1997mathematical . This reformulation removes the need for ODE solvers in backpropagation, improving computational efficiency while preserving accuracy. By learning accelerations directly, the model captures the instantaneous dynamics of the system, leading to better generalization and more stable long-term predictions. Moreover, FNODE naturally accommodates various numerical differentiation schemes for computing acceleration targets from position data, including finite differences and spectral methods, allowing practitioners to balance between noise robustness and accuracy based on their data characteristics fornberg1988generation .
@@
Directly learning accelerations requires accurate derivative information, which can be difficult to obtain from noisy trajectory data. To address this, we use a hybrid scheme that combines the accuracy of Fast Fourier Transform (FFT)-based spectral differentiation trefethen2000spectral with the stability of finite differences (FD) near trajectory boundaries. Our numerical experiments show that, with proper differentiation techniques, the advantages of acceleration-based learning outweigh the challenges of derivative estimation.
This paper makes the following contributions:
@@
    We evaluate FNODE on a diverse set of multibody dynamics benchmarks, showing consistent improvements over baselines—including MBDNODEs, LSTMs, and FCNNs—in long-term prediction accuracy, out-of-distribution generalization, and computational efficiency.
@@ 2.1 The Multibody Dynamics Problem
We employ a MBD formulation in redundant generalized coordinates wherein Lagrange multipliers are used to incorporate constraint equations directly into the equations of motion shabana2020dynamics .
@@
Above, the mass matrix $\mathbf{M}$ encodes the system’s inertia, while $\mathbf{\Phi_{q}}$ is the Jacobian of the kinematic constraints. The generalized coordinates are collected in $\mathbf{q}$, with $\ddot{\mathbf{q}}$ denoting accelerations. The Lagrange multipliers $\lambda$ correspond to the constraint forces. The right-hand side includes $\mathbf{F_{e}}$, the vector of external and velocity-dependent forces, and $\mathbf{\gamma_{c}}$, the contribution from the second time derivative of the constraints.
@@ 2.2.1 Neural Ordinary Differential Equation (NODE)
2.2.2 Flow-Matching in Neural Ordinary Differential Equation
FNODE operates on an augmented state space defined as $\bm{Z}(t)=(\bm{z}(t),\dot{\bm{z}}(t))^{T}$ that is used to model the acceleration directly:
@@ 2.2.2 Flow-Matching in Neural Ordinary Differential Equation
$\bm{z}(t,\bm{\mu})=(\bm{z}^{1}(t,\bm{\mu}),...,\bm{z}^{n_{z}}(t,\bm{\mu}))^{T}\in\mathbb{R}^{n_{z}},$ (6)
$\dot{\bm{z}}(t,\bm{\mu})=(\dot{\bm{z}}^{1}(t,\bm{\mu}),\dots,\dot{\bm{z}}^{n_{z}}(t,\bm{\mu}))^{T}\in\mathbb{R}^{n_{z}}\;.$ (7)
@@
$f:\mathbb{R}^{2n_{z}}\times\mathbb{R}^{n_{\bm{\mu}}}\to\mathbb{R}^{n_{z}}\;.$ (8)
@@
Figure 1: The discretized forward pass for FNODE for general MBD
Figure 1 shows the discretized forward pass of FNODE. The network maps the system state to accelerations using the three-layer architecture in Table 1, with Tanh/ReLU activations and Xavier/Kaiming initialization.
Table 1: MBD-NODE Architecture
@@
Input Layer $2n_{z}$ [Tanh,ReLU] [Xavier, Kaiming]
Hidden Layer 1 $d_{\text{width}}$ [Tanh,ReLU] [Xavier, Kaiming]
Hidden Layer 2 $d_{\text{width}}$ [Tanh,ReLU] [Xavier, Kaiming]
Output Layer $n_{z}$ - [Xavier, Kaiming]
@@ 2.3 Calculation of Acceleration
Two primary approaches are employed depending on the characteristics of the available dataset: Fast Fourier Transform (FFT) spectral differentiation and Finite Difference (FD) methods.
@@ 2.3.1 FFT Spectral Differentiation
For dynamic systems which satisfies periodicity assumptions, FFT-based spectral differentiation provides is known for high accuracy in time gradient calculation. Given a periodic function $z(t)$ with time interval $L=N\Delta t$ and uniform time step $\Delta t$, the Fourier series representation is:
@@
The time gradient computed by the spectral representation assumes then the expression:
@@
Due to aliasing errors and Gibbs phenomenon, the results may contain spurious high-frequency oscillations. The Gibbs error arises from discontinuity in periodical expansion:
$\Delta_{boundary}=z(N-1)-z(0)\neq 0\;.$ (12)
@@
This creates a constant-amplitude error across all high frequency components as shown in Eq. (14), leading to a total Gibbs error
$\epsilon_{Gibbs}\approx\frac{|\Delta_{boundary}|}{2\Delta t}\;.$ (15)
The error is independent of $N$ but inversely proportional to the time step, which means it would be particularly problematic for fine temporal resolution. To solve the Gibbs phenomenon problem described in Eq. (15), least-square detrending is applied
@@
where $\ell(t)$ is the linear trend function to approximate the trend of the raw dataset. The detrended dataset is formulated using Eq.(16) as
@@
Even though the detrend narrows the gap of the two ends of dataset as described in Eq. (18), the first order derivative of the dataset is still not continuous. In order to further reduce Gibbs error while maintain the spectrum information as mush as possible, we adopt mirror reflection and cosine taper.
@@
The complete FFT-based differentiation process is formulated as follows. First, we compute the discrete Fourier transform (DFT) of the preprocessed signal $\mathbf{z}_{e}[n]$ with the applied window function $w[n]$ from Eq. (20):
@@
This approach combines spectral differentiation with noise suppression, yielding smooth and accurate derivative estimates for FNODE training.
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Using Taylor expansions, the first derivative can be approximated by atkinson2008introduction :
$z^{\prime}(t)\approx\frac{z(t+\Delta t)-z(t-\Delta t)}{2\Delta t}\quad\text{(central, second order)}\;,$ (23a)
with forward and backward differences used at boundaries:
$z^{\prime}(t)\approx\frac{z(t+\Delta t)-z(t)}{\Delta t}\;,\quad z^{\prime}(t)\approx\frac{z(t)-z(t-\Delta t)}{\Delta t}\;.$ (23b)
The central scheme has truncation error $\mathcal{O}(\Delta t^{2})$, while forward and backward schemes are $\mathcal{O}(\Delta t)$. Although less accurate than spectral differentiation, FD methods are more stable for short trajectories and at dataset boundaries, making them complementary in our hybrid FFT-FD approach.
2.4 Error Propagation in Integration
While FNODE avoids numerical integration during training, its long-term accuracy depends on how errors propagate through the ODE solver. To carry out a basic error analysis following a classical framework hairer1993classical ; brenan1996numerical , let the true system state at step $n$ be $Z_{n}$, advanced by the exact evolution operator $\Phi_{g}$, and let $\hat{Z}_{n}$ denote the predicted state advanced by the learned solver $\Phi_{f}$. The global error is $E_{n}=\hat{Z}_{n}-Z_{n}$. The exact and predicted updates are
$Z_{n+1}=\Phi_{g}(Z_{n}),\qquad\hat{Z}_{n+1}=\Phi_{f}(\hat{Z}_{n}),$

### v3

@@ Abstract
Data-driven modeling of constrained multibody dynamics remains challenged by (i) the training cost of Neural ODEs, which typically require backpropagation through an ODE solver, and (ii) error accumulation in rollout predictions. We introduce a Flow-Matching Neural ODE (FNODE) framework that learns the acceleration mapping directly from trajectory data by supervising accelerations rather than integrated states, turning training into a supervised regression problem and eliminating the ODE-adjoint/solver backpropagation bottleneck. Acceleration targets are obtained efficiently via numerical differentiation using a hybrid fast Fourier transform (FFT) and finite-difference (FD) scheme. Kinematic constraints are enforced through coordinate partitioning: FNODE learns accelerations only for the independent generalized coordinates, while the dependent coordinates are recovered by solving the position-level constraint equations. We evaluate FNODE on single and triple mass–spring–damper systems, a double pendulum, a slider–crank with and without friction, a vehicle model, and a cart–pole, and compare against MBD-NODE, LSTM, and fully connected baselines. Across these benchmarks, FNODE achieves improved prediction accuracy and training/runtime efficiency, while maintaining constraint satisfaction through the partitioning procedure. Our code and scripts are released as open source to support reproducibility and follow-on research.
Keywords Multibody dynamics $\cdot$ Neural ODE $\cdot$ Constrained dynamics $\cdot$ Scientific machine learning
@@ 1 Introduction
Multibody dynamics is relevant in simulation-based engineering, with applications in numerous fields, e.g., robotics [1], vehicle and train dynamics [2, 3], biomechanics [4], aircraft dynamics [5] and molecular dynamics [6], to name a few. Traditionally, modeling multibody systems has relied on first-principles approaches such as Newton-Euler formulations [7], Lagrangian mechanics [8], and Hamiltonian mechanics [9]. While these classical methods provide accurate and interpretable models, they are challenged in scenarios where the underlying dynamics are partially unknown, highly complex, or subject to unmodeled effects such as friction, contact forces, or material nonlinearities [10].
@@
Machine learning has opened the door to data-driven simulation of dynamical systems, offering the potential to learn complex behaviors directly from observational data [11, 12]. Early approaches in this domain employed classical machine learning architectures such as LSTM [13, 14] and FCNN [15] to learn discrete-time mappings from current states to future states. In addition, recent work [16, 17] explored fixed-time increment DNN methods combined with dimensionality reduction techniques for efficient multibody dynamics simulation. While these methods have shown promise in short-term prediction tasks, they demonstrated limited out-of-distribution generalization. Specifically, these black-box predictors learn fixed-timestep transitions $f:s_{t}\mapsto s_{t+\Delta t}$, which limits their temporal resolution and also causes error accumulation that grows exponentially with the prediction horizon [18]. Furthermore, such discrete-time approaches often fail to enforce fundamental conservation laws when applied in conjunction with conservative systems [19]. Physics-informed approaches have been proposed to address these limitations by incorporating physical constraints and laws into the model, presenting improved reliability through enforcement of physical principles in neural network training [20, 21, 22]. Neural Ordinary Differential Equations (NODEs) [23] marked a major step in continuous-time learning by representing system dynamics as neural vector fields integrated with standard ODE solvers. This formulation handles irregular time sampling and directly connects learning with established numerical integration theory [24, 25].
Meanwhile, hybrid methods combining first-principle models with NODEs have shown promise in specific applications through automatic parameter tuning [26]. However, these approaches typically assume that the functional form of the governing equations (or physics structure) is known a priori, which can limit their ability to generalize when the underlying mechanisms are partially unknown or misspecified. To improve physical consistency, several structure-preserving learning directions have been proposed, including energy-based formulations such as Hamiltonian and Lagrangian Neural Networks [27, 28], symplectic-geometry-based approaches such as Symplectic ODE-Net and Symplectic Machine Learning (Sym-ML) frameworks [29, 30], and methods that target numerical stability and explicit constraint satisfaction, e.g., Stabilized Neural Differential Equations [31]. For flow-matching-based methods, recent work proposes a contrastive feature-alignment remedy for systematic failure modes that arise in the low-noise regime [32]. Separately, another work develops improved training strategies for rectified flows [33]. In a different modeling paradigm, operator-learning methods such as DeepONets [34] and Fourier Neural Operators [35] learn solution operators to generalize across parameterized families of dynamical systems. However, despite their strengths, NODEs suffer from a bottleneck: the adjoint sensitivity method for backpropagation requires solving an additional backward ODE, which increases computational cost and memory use, and limits scalability for stiff or high-dimensional systems [36, 37].
@@
Recent advances in flow-matching algorithms [38, 39] show that vector fields can be learned by directly supervising derivatives at sampled points, avoiding backward ODE solves [40]. The approach of learning state-acceleration mapping relationship has been successfully applied to simulate flexible body dynamics [41]. Applying these ideas to constrained multibody dynamics, we train neural networks to approximate accelerations directly, eliminating costly integration in the learning loop. We propose FNODE, a framework for data-driven modeling of multibody systems. Instead of predicting future states via time integration, FNODE trains a neural network to learn directly the acceleration vector field – the natural quantity in mechanical systems governed by force laws [42]. Kinematic constraints are enforced through a coordinate-partitioning approach [43], in which FNODE learns the acceleration field only for the independent generalized coordinates, while the dependent coordinates are recovered implicitly by enforcing the kinematic constraint equations at the position level. This reformulation removes the need for ODE solvers in backpropagation, improving computational efficiency while preserving accuracy. By learning accelerations directly, the model captures the instantaneous dynamics of the system, leading to better generalization and more stable long-term predictions. Moreover, FNODE naturally accommodates various numerical differentiation schemes for computing acceleration targets from position data, including finite differences and FFT-based spectral methods, allowing practitioners to balance between noise robustness and accuracy based on their data characteristics [44, 45].
@@
Directly learning accelerations requires accurate derivative information, which can be difficult to obtain from noisy trajectory data. To address this, we use a hybrid scheme that combines the accuracy of FFT-based spectral differentiation with the stability of FD near trajectory boundaries. Our numerical experiments show that, with proper differentiation techniques, the advantages of acceleration-based learning outweigh the challenges of derivative estimation.
The following are our main contributions in this paper:
@@
    We evaluate FNODE on a diverse set of constrained multibody dynamics benchmarks, showing consistent improvements over baselines—including MBD-NODE, LSTMs, and FCNNs—in long-term prediction accuracy, out-of-distribution generalization, and computational efficiency.
@@ 2.1 The Multibody Dynamics Problem
We employ an MBD formulation in redundant generalized coordinates in which Lagrange multipliers are used to account for the presence of kinematic constraints in the equations of motion [46].
@@
Above, the mass matrix $\mathbf{M}$ encodes the system’s inertia, while $\mathbf{\Phi_{q}}$ is the Jacobian of the kinematic constraints. The generalized coordinates are collected in $\mathbf{q}$, with $\ddot{\mathbf{q}}$ denoting accelerations. The Lagrange multipliers $\lambda$ correspond to the constraint forces. The right-hand side includes $\mathbf{F_{e}}$, the vector of external and velocity-dependent forces, and $\mathbf{\gamma_{c}}$, the contribution from the second time derivative of the kinematic constraints. For details, see [46, 47].
@@ 2.2.1 Neural Ordinary Differential Equation (NODE)
2.2.2 Flow-Matching in NODE
FNODE operates on an augmented state space defined as $\bm{Z}(t)=(\bm{z}^{T}(t),\dot{\bm{z}}^{T}(t))^{T}$ that is used to model the acceleration directly:
@@ 2.2.2 Flow-Matching in NODE
$\bm{z}(t,\bm{\mu})=(z^{1}(t,\bm{\mu}),...,z^{n_{z}}(t,\bm{\mu}))^{T}\in\mathbb{R}^{n_{z}},$ (6)
$\dot{\bm{z}}(t,\bm{\mu})=(\dot{z}^{1}(t,\bm{\mu}),\dots,\dot{z}^{n_{z}}(t,\bm{\mu}))^{T}\in\mathbb{R}^{n_{z}}\;.$ (7)
@@
$f(\mathbf{z}(t),t;\mathbf{\Theta}):\mathbb{R}^{2n_{z}}\times\mathbb{R}^{n_{\bm{\mu}}}\to\mathbb{R}^{n_{z}}\;.$ (8)
@@
Figure 1: Discretized forward pass of the FNODE framework for a general multibody dynamics (MBD) system. The solid arrows represent the data flow during both training and inference. The dashed arrows indicate latent inference operations. The dotted arrows denote the calculation of time derivative (training target). During training, “ground-truth” accelerations are computed from trajectory data using FFT- or finite-difference (FD)-based differentiation. Once trained, the neural model $f(\cdot;\bm{\Theta})$ predicts accelerations that are supplied to an ODE solver during inference to generate the system’s time evolution.
Figure 1 shows the discretized forward pass of FNODE. The network maps the system state to accelerations using the three-layer architecture in Table 1, with tanh activations and Xavier initialization.
Table 1: FNODE Architecture
@@
Input Layer $2n_{z}$ tanh Xavier
Hidden Layer 1 $d_{\text{width}}$ tanh Xavier
Hidden Layer 2 $d_{\text{width}}$ tanh Xavier
Output Layer $n_{z}$ - Xavier
@@ 2.3 Calculation of Acceleration
Two primary approaches are employed depending on the characteristics of the available dataset: FFT spectral differentiation and FD methods.
@@ 2.3.1 FFT Spectral Differentiation
Continuous Fourier Representation
For dynamic systems which satisfy periodicity assumptions, FFT-based spectral differentiation provides high accuracy in time derivative calculation. We begin by presenting the ideal mathematical framework for a smooth periodic function $z(t)$ with time interval $L=N\Delta t$ and uniform time step $\Delta t$. The exact Fourier series representation is:
@@
The time derivative computed by the exact spectral representation assumes then the expression:
@@
Discrete Implementation and Artifacts
Eqs. (10)–(11) describe the ideal continuous-time Fourier representation for a smooth periodic trajectory, and motivate spectral differentiation through a frequency-domain multiplier. In practice, however, we only observe a finite set of samples and compute derivatives using the discrete Fourier transform (DFT)/FFT.
In this discrete setting, sampling (finite $\Delta t$) introduces a Nyquist limit $f_{N}=\tfrac{1}{2\Delta t}$ and thus possible aliasing of unresolved high-frequency content, while finite-length windowing and imperfect periodicity introduce spectral leakage. These artifacts are not properties of the exact Fourier series itself, but become central once differentiation is implemented on sampled data. Moreover, since differentiation multiplies each mode by its (angular) frequency, high-frequency components—including aliased noise—are amplified in the derivative estimates. Accordingly, in the sampled DFT/FFT implementation, aliasing and finite-window effects (including Gibbs-type ringing) may introduce spurious high-frequency oscillations.
One particularly important manifestation of these discrete-sampling artifacts is the Gibbs error, which arises from boundary discontinuity in the periodic expansion:
$\Delta_{boundary}=z((N-1)\Delta t)-z(0)\neq 0\;.$ (12)
@@
This implies that, in the presence of a boundary discontinuity, the differentiated Fourier modes do not decay with $|k|$ in the high-frequency range; instead, their magnitude remains $O\!\left(\tfrac{|\Delta_{\text{boundary}}|}{\Delta t}\right)$ as $|k|\to\infty$ (cf. Eq. (14)). As a result, the truncated Fourier reconstruction of the derivative can exhibit spurious high-frequency oscillations (Gibbs-type ringing), with a characteristic (order-of-magnitude) error scale
$\epsilon_{\text{Gibbs}}\sim\frac{|\Delta_{\text{boundary}}|}{2\Delta t}\;.$ (15)
The error is independent of $N$ but inversely proportional to the time step, which means it would be particularly problematic for fine temporal resolution.
To address the Gibbs phenomenon problem described in Eq. (15) and other boundary discontinuities, least-square detrending is applied:
@@
where $\ell(t)$ is the linear trend function to approximate the trend of the raw dataset. The detrended dataset is formulated using Eq. 16 as
@@
Even though the detrend narrows the gap of the two ends of dataset as described in Eq. (18), the first-order derivative of the dataset is still not continuous. In order to further reduce Gibbs error while maintain the spectrum information as much as possible, we adopt mirror reflection and cosine taper.
@@
Complete FFT-Based Differentiation Pipeline
The complete FFT-based differentiation process is formulated as follows. Let $t_{n}=n\Delta t$ for $n=0,1,\dots,N-1$, and denote the sampled (preprocessed) sequence by $\mathbf{z}_{e}[n]:=\mathbf{z}_{e}(t_{n})$. First, we compute the discrete Fourier transform (DFT) of the preprocessed signal $\mathbf{z}_{e}[n]$ with the applied window function $w[n]$ from Eq. (20):
@@
This approach combines spectral differentiation with noise suppression. The FFT-based pipeline is most reliable for smooth, near-periodic trajectories after preprocessing; for non-periodic or non-smooth signals (e.g., impacts/contact events), Fourier truncation can lead to Gibbs oscillations and boundary artifacts, and any explicit differentiation operator (FFT or FD) amplifies measurement noise. Therefore, when FNODE is trained using explicit acceleration labels obtained via differentiation, the current differentiation choices primarily limit applicability to smooth regimes. Note that practical error diagnostics for differentiated signals, including a reconstruction-consistency check based on integrating the derivatives and comparing against the original trajectory, are described in Section 2.3.2.
Since acceleration labels are obtained via explicit differentiation, occasional high-frequency artifacts (e.g., due to boundary treatment, FFT/finite-difference effects, or nonsmooth events such as frictional transitions) can manifest as outliers in the target residual. In such cases, replacing the mean-squared error (MSE) with a robust regression loss can reduce sensitivity to outliers without changing the FNODE architecture. A common choice is the Huber loss applied to the component-wise residual $r=\hat{a}-a$:
$\rho_{\delta}(r)=\begin{cases}\tfrac{1}{2}r^{2},&|r|\leq\delta,\\[4.0pt] \delta\left(|r|-\tfrac{1}{2}\delta\right),&|r|>\delta.\end{cases}$ (23)
For small residuals the loss behaves like MSE, while large residuals are down-weighted (linear growth), making training less sensitive to impulsive artifacts. In this manuscript we focus on noise-free trajectories for clarity and do not perform a systematic study over injected noise types/levels (e.g., Gaussian or impulse noise); however, robust losses such as (23) are a standard mitigation when differentiated targets contain occasional spikes.
@@ 2.3.2 Finite Difference Method
For non-periodic datasets, finite-difference (FD) schemes provide a simple and robust alternative to FFT-based differentiation. Let $z^{\prime}_{t}$ denote the FD estimate of $z^{\prime}(t)$. Using the central-difference formula, $z^{\prime}_{t}=\tfrac{z(t+\Delta t)-z(t-\Delta t)}{2\Delta t}$, and thus $z^{\prime}(t)=z^{\prime}_{t}+\mathcal{O}(\Delta t^{2})$ [48]. At the boundaries, we use one-sided formulas, $z^{\prime}_{t}=\tfrac{z(t+\Delta t)-z(t)}{\Delta t}$ and $z^{\prime}_{t}=\tfrac{z(t)-z(t-\Delta t)}{\Delta t}$, and hence $z^{\prime}(t)=z^{\prime}_{t}+\mathcal{O}(\Delta t)$ in either case. Although less accurate than spectral differentiation, FD methods are more stable for short trajectories and at dataset boundaries, making them complementary in our hybrid FFT–FD approach.
Higher-order FD discretizations are also available. For example, let $z^{\prime(5)}_{t}$ denote the five-point central FD estimate,
$z^{\prime(5)}_{t}=\tfrac{-z(t+2\Delta t)+8z(t+\Delta t)-8z(t-\Delta t)+z(t-2\Delta t)}{12\Delta t}\;,$
so that $z^{\prime}(t)=z^{\prime(5)}_{t}+\mathcal{O}(\Delta t^{4})$, together with matching higher-order one-sided stencils near boundaries. In our setting, trajectories are short and derivative labels are obtained from sampled data; wider stencils increase the number of samples affected by boundary treatment and can amplify high-frequency noise. We therefore adopt the three-point central stencil in the interior and one-sided stencils at the endpoints as a conservative accuracy–robustness tradeoff.
To obtain a practical estimate of differentiation error on a given dataset, one may use two complementary consistency checks. First, compute derivative estimates with two stencil orders and define a relative disagreement indicator $\eta^{\prime}=\tfrac{\|z^{\prime(5)}-z^{\prime(3)}\|_{2}}{\|z^{\prime(5)}\|_{2}}$, where $z^{\prime(3)}$ and $z^{\prime(5)}$ denote the three-point and five-point FD estimates of $z^{\prime}(t)$ (and analogously for $z^{\prime\prime}$), and $\|\cdot\|_{2}$ denotes the discrete $\ell_{2}$ norm over the sampled trajectory. Second, define a reconstruction-consistency indicator by integrating the differentiated signals and comparing to the original trajectory, e.g., $\tilde{z}_{n+1}=\tilde{z}_{n}+\Delta t\,z^{\prime}_{n}$ with $\tilde{z}_{0}=z_{0}$, and form the normalized reconstruction residual $\eta_{z}=\tfrac{\|z-\tilde{z}\|_{2}}{\|z\|_{2}}$. These diagnostics provide a practical way to quantify sensitivity to the differentiation scheme and to detect differentiation artifacts when constructing acceleration targets for FNODE training; a systematic evaluation of $\eta^{\prime}$ and $\eta_{z}$ across datasets is deferred to future work and falls outside the scope of the present contribution.
2.4 Qualitative Analysis of Error Propagation in Integration
While FNODE avoids numerical integration during training, its long-term accuracy depends on how errors propagate through the ODE solver. To carry out a basic error analysis following a classical framework [49, 50], let the true system state at step $n$ be $Z_{n}$, advanced by the exact evolution operator $\Phi$, and let $\hat{Z}_{n}$ denote the predicted state advanced by the learned solver $\Phi_{\Theta}$. The global error is $E_{n}=\hat{Z}_{n}-Z_{n}$. The exact and predicted updates are
$Z_{n+1}=\Phi(Z_{n}),\qquad\hat{Z}_{n+1}=\Phi_{\Theta}(\hat{Z}_{n}),$